# Another look at inference after prediction

### v2

@@ B.1 Asymptotic expansion of $\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$
Theorem S1 (Influence function expansion of $\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$)
@@
Proof  The only difference in studying the asymptotic properties of $\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$ compared to $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ is that we rely on the asymptotic linear expansion of $\hat{\bm{\gamma}}^{\text{\tiny{unlab}}}$ rather than $\hat{\bm{\gamma}}^{\text{\tiny{all}}}$. Using the same argument as in the proof of Theorem 1,
@@ B.2 Comparison of $\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$ $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$
Corollary S1.1 (Comparison of $\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$ with $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$)
@@
$\Delta_{\text{aVar}}(\hat{\bm{\beta}}^{\text{\tiny{lab}}},\hat{\bm{\beta}}^{\text{\tiny{PPI}}})$ $=\frac{1}{\pi}{\bf{A}}\left(2\boldsymbol{\Sigma}_{Y,\widehat{Y}}-\frac{1}{1-\pi}\boldsymbol{\Sigma}_{\widehat{Y},\widehat{Y}}\right){\bf{A}}.$
$\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$ is therefore more efficient than $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ if and only if $2\boldsymbol{\Sigma}_{Y,\widehat{Y}}\succ\frac{1}{1-\pi}\boldsymbol{\Sigma}_{\widehat{Y},\widehat{Y}}$. A sufficient condition for $\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$ to have lower variance than $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ is that
@@
Proof  From the proof for Corollary 1.1, $n[\text{aVar}(\hat{\bm{\beta}}^{\text{\tiny{lab}}})]=\frac{1}{\pi}{\bf{A}}\boldsymbol{\Sigma}_{Y,Y}{\bf{A}}.$ By Theorem S1,
@@
$=\frac{1}{\pi}{\bf{A}}\boldsymbol{\Sigma}_{Y,Y}{\bf{A}}+\frac{1}{\pi(1-\pi)}{\bf{A}}\boldsymbol{\Sigma}_{\widehat{Y},\widehat{Y}}{\bf{A}}-\frac{2}{\pi}{\bf{A}}\boldsymbol{\Sigma}_{Y,\widehat{Y}}{\bf{A}}$
$=\frac{1}{\pi}{\bf{A}}\left[\boldsymbol{\Sigma}_{Y,Y}+\frac{1}{1-\pi}\boldsymbol{\Sigma}_{\widehat{Y},\widehat{Y}}-2\boldsymbol{\Sigma}_{Y,\widehat{Y}}\right]{\bf{A}}.$
@@
$\Delta_{\text{aVar}}(\hat{\bm{\beta}}^{\text{\tiny{lab}}},\hat{\bm{\beta}}^{\text{\tiny{PPI}}})=\frac{1}{\pi}{\bf{A}}\left[2\boldsymbol{\Sigma}_{Y,\widehat{Y}}-\frac{1}{1-\pi}\boldsymbol{\Sigma}_{\widehat{Y},\widehat{Y}}\right]{\bf{A}}.$
$\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$ is therefore more efficient than $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ if and only if $2\boldsymbol{\Sigma}_{Y,\widehat{Y}}\succ\frac{1}{1-\pi}\boldsymbol{\Sigma}_{\widehat{Y},\widehat{Y}}$. Note that if $2\mathbb{E}\left(\epsilon_{Y}\epsilon_{\widehat{Y}}\mid{\bf{X}}\right)>\frac{1}{1-\pi}\mathbb{E}\left(\epsilon_{\widehat{Y}}^{2}\mid{\bf{X}}\right)$ or equivalently if $\mathbb{E}\left[\left(\epsilon_{\widehat{Y}}-\epsilon_{Y}\right)^{2}\mid{\bf{X}}\right]<\mathbb{E}\left(\epsilon_{Y}^{2}\mid{\bf{X}}\right)+\left(1-\frac{1}{1-\pi}\right)\mathbb{E}\left(\epsilon_{\widehat{Y}}^{2}\mid{\bf{X}}\right),$ the diagonal entries of $\Delta_{\text{aVar}}(\hat{\bm{\beta}}^{\text{\tiny{lab}}},\hat{\bm{\beta}}^{\text{\tiny{PPI}}})$ are positive and $\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$ will have a lower variance than $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$.
@@ B.3 Comparison of $\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$ $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$
$2\boldsymbol{\Sigma}_{Y,\widehat{Y}}\succ\frac{1}{1-\pi}\boldsymbol{\Sigma}_{\widehat{Y},\widehat{Y}}$
@@
$2\boldsymbol{\Sigma}_{Y,\widehat{Y}}\succ\boldsymbol{\Sigma}_{\widehat{Y},\widehat{Y}}.$
@@
$=\left(\frac{1}{\pi}-1\right){\bf{A}}\left(2\boldsymbol{\Sigma}_{Y,\widehat{Y}}-\boldsymbol{\Sigma}_{\widehat{Y},\widehat{Y}}\right){\bf{A}}-\frac{1}{\pi}{\bf{A}}\left(2\boldsymbol{\Sigma}_{Y,\widehat{Y}}-\frac{1}{1-\pi}\boldsymbol{\Sigma}_{\widehat{Y},\widehat{Y}}\right){\bf{A}}$
$={\bf{A}}\left(\frac{2-\pi}{1-\pi}\boldsymbol{\Sigma}_{\widehat{Y},\widehat{Y}}-2\boldsymbol{\Sigma}_{Y,\widehat{Y}}\right){\bf{A}}.$
Note that if $\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$ is less efficient than $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ then $\frac{1}{1-\pi}\boldsymbol{\Sigma}_{\widehat{Y},\widehat{Y}}\succ 2\boldsymbol{\Sigma}_{Y,\widehat{Y}}$, implying
$\frac{2-\pi}{1-\pi}\boldsymbol{\Sigma}_{\widehat{Y},\widehat{Y}}\succ 2\boldsymbol{\Sigma}_{Y,\widehat{Y}}.$
@@ C.1 Proof of Theorem 2
Theorem 2 (Influence function expansion of $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$)
Following Chen and Chen (2000),
@@
Proof  The proof follows from Chen and Chen (2000), but is detailed for completeness. Recall that
$\begin{split}\widehat{{\bf{W}}}^{\text{\tiny{CC}}}&=\widehat{{\bf{A}}}\left\{\widehat{\text{Cov}}\left[\phi(Y,{\bf{X}};\bm{\beta}^{*}),\phi(\widehat{Y},{\bf{X}};\bm{\gamma}^{*})\right]\right\}\{\widehat{\text{Var}}[\phi(\widehat{Y},{\bf{X}};\bm{\gamma}^{*})]\}^{-1}\widehat{{\bf{A}}}^{-1}\\ &=\widehat{{\bf{A}}}\widehat{\boldsymbol{\Sigma}}_{Y,\widehat{Y}}\widehat{\boldsymbol{\Sigma}}^{-1}_{\widehat{Y},\widehat{Y}}\widehat{{\bf{A}}}^{-1}\end{split}$
@@
$\begin{split}&\widehat{{\bf{A}}}=n_{\text{\tiny{lab}}}\left(\sum_{i=1}^{n_{\text{\tiny{lab}}}}{\bf{X}}_{i}{\bf{X}}_{i}^{{\sf\scriptscriptstyle{T}}}\right)^{-1},\\ &\widehat{\boldsymbol{\Sigma}}_{Y,\widehat{Y}}=n_{\text{\tiny{lab}}}^{-1}\sum_{i=1}^{n_{\text{\tiny{lab}}}}\left(Y_{i}-{\bf{X}}_{i}^{{\sf\scriptscriptstyle{T}}}\hat{\bm{\beta}}^{\text{\tiny{lab}}}\right)\left(\hat{Y}_{i}-{\bf{X}}_{i}^{{\sf\scriptscriptstyle{T}}}\hat{\bm{\gamma}}^{\text{\tiny{lab}}}\right){\bf{X}}_{i}{\bf{X}}_{i}^{{\sf\scriptscriptstyle{T}}},\\ &\mbox{and }\widehat{\boldsymbol{\Sigma}}_{\widehat{Y},\widehat{Y}}=n_{\text{\tiny{lab}}}^{-1}\sum_{i=1}^{n_{\text{\tiny{lab}}}}\left(\hat{Y}_{i}-{\bf{X}}_{i}^{{\sf\scriptscriptstyle{T}}}\hat{\bm{\gamma}}^{\text{\tiny{lab}}}\right)^{2}{\bf{X}}_{i}{\bf{X}}_{i}^{{\sf\scriptscriptstyle{T}}}.\end{split}$
@@
$n^{\frac{1}{2}}\left[\widehat{{\bf{W}}}^{\text{\tiny{CC}}}(\hat{\bm{\gamma}}^{\text{\tiny{lab}}}-\bm{\gamma}^{*})\right]={\bf{A}}\boldsymbol{\Sigma}_{Y,\widehat{Y}}\boldsymbol{\Sigma}^{-1}_{\widehat{Y},\widehat{Y}}\left\{n^{-\frac{1}{2}}\sum_{i=1}^{n}\left[{\bf{X}}_{i}\left(\widehat{Y}_{i}-{\bf{X}}_{i}^{{\sf\scriptscriptstyle{T}}}\bm{\gamma}^{*}\right)\right]\frac{R_{i}}{\pi}\right\}+o_{p}(1)$
@@
$n^{\frac{1}{2}}\left[\widehat{{\bf{W}}}^{\text{\tiny{CC}}}(\hat{\bm{\gamma}}^{\text{\tiny{all}}}-\bm{\gamma}^{*})\right]={\bf{A}}\boldsymbol{\Sigma}_{Y,\widehat{Y}}\boldsymbol{\Sigma}^{-1}_{\widehat{Y},\widehat{Y}}\left\{n^{-\frac{1}{2}}\sum_{i=1}^{n}\left[{\bf{X}}_{i}\left(\widehat{Y}_{i}-{\bf{X}}_{i}^{{\sf\scriptscriptstyle{T}}}\bm{\gamma}^{*}\right)\right]\right\}+o_{p}(1).$
@@
Corollary 2.1 ($\hat{\bm{\beta}}^{\text{\tiny{CC}}}$ is at least as efficient as $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$)
@@
$\Delta_{\text{aVar}}(\hat{\bm{\beta}}^{\text{\tiny{lab}}},\hat{\bm{\beta}}^{\text{\tiny{CC}}})=\left(\frac{1}{\pi}-1\right){\bf{A}}\boldsymbol{\Sigma}_{Y,\widehat{Y}}\boldsymbol{\Sigma}^{-1}_{\widehat{Y},\widehat{Y}}\boldsymbol{\Sigma}_{Y,\widehat{Y}}{\bf{A}}\succeq{\bf{0}}.$
@@
Proof  From the proof for Corollary 1.1, $n[\text{aVar}(\hat{\bm{\beta}}^{\text{\tiny{lab}}})]=\frac{1}{\pi}{\bf{A}}\boldsymbol{\Sigma}_{Y,Y}{\bf{A}}.$ By Theorem 2.1, $n[\text{aVar}(\hat{\bm{\beta}}^{\text{\tiny{CC}}})]$ is given by
$\begin{split}&\mathbb{E}\left\{\left[{\bf{A}}\bm{\phi}(Y,{\bf{X}};\bm{\beta}^{*})\right]^{\otimes 2}\frac{R^{2}}{\pi^{2}}\right\}+\mathbb{E}\left\{\left[{\bf{A}}\boldsymbol{\Sigma}_{Y,\widehat{Y}}\boldsymbol{\Sigma}^{-1}_{\widehat{Y},\widehat{Y}}\bm{\phi}(\widehat{Y},{\bf{X}};\bm{\gamma}^{*})\right]^{\otimes 2}\frac{(\pi-R)^{2}}{\pi^{2}}\right\}\\ +&2\mathbb{E}\left\{\left[{\bf{A}}\bm{\phi}(Y,{\bf{X}};\bm{\beta}^{*})\right]\left[{\bf{A}}\boldsymbol{\Sigma}_{Y,\widehat{Y}}\boldsymbol{\Sigma}^{-1}_{\widehat{Y},\widehat{Y}}\bm{\phi}(\widehat{Y},{\bf{X}};\bm{\gamma}^{*})\right]^{{\sf\scriptscriptstyle{T}}}\frac{R(\pi-R)}{\pi^{2}}\right\}\\ =&\frac{1}{\pi}{\bf{A}}\boldsymbol{\Sigma}_{Y,Y}{\bf{A}}+\frac{1-\pi}{\pi}{\bf{A}}\boldsymbol{\Sigma}_{Y,\widehat{Y}}\boldsymbol{\Sigma}^{-1}_{\widehat{Y},\widehat{Y}}\boldsymbol{\Sigma}_{Y,\widehat{Y}}{\bf{A}}-2\frac{1-\pi}{\pi}{\bf{A}}\boldsymbol{\Sigma}_{Y,\widehat{Y}}\boldsymbol{\Sigma}^{-1}_{\widehat{Y},\widehat{Y}}\boldsymbol{\Sigma}_{Y,\widehat{Y}}{\bf{A}}\\ =&\frac{1}{\pi}{\bf{A}}\boldsymbol{\Sigma}_{Y,Y}{\bf{A}}-\frac{1-\pi}{\pi}{\bf{A}}\boldsymbol{\Sigma}_{Y,\widehat{Y}}\boldsymbol{\Sigma}^{-1}_{\widehat{Y},\widehat{Y}}\boldsymbol{\Sigma}_{Y,\widehat{Y}}{\bf{A}}.\end{split}$
Hence, $\Delta_{\text{aVar}}(\hat{\bm{\beta}}^{\text{\tiny{lab}}},\hat{\bm{\beta}}^{\text{\tiny{CC}}})=\frac{1-\pi}{\pi}{\bf{A}}\boldsymbol{\Sigma}_{Y,\widehat{Y}}\boldsymbol{\Sigma}^{-1}_{\widehat{Y},\widehat{Y}}\boldsymbol{\Sigma}_{Y,\widehat{Y}}{\bf{A}}\succeq 0$.
@@
Corollary 2.2 ($\hat{\bm{\beta}}^{\text{\tiny{CC}}}$ is the most efficient estimator in its class)
Let the mean squared error (MSE) of $\hat{\bm{\beta}}$ for estimating $\bm{\beta}^{*}$ be defined as $\mathbb{E}[(\hat{\bm{\beta}}-\bm{\beta}^{*})^{{\sf\scriptscriptstyle{T}}}(\hat{\bm{\beta}}-\bm{\beta}^{*})]$. Analogously to the well-known result for the best linear predictor (e.g., Bickel and Doksum (2015), page 40), ${\bf{W}}^{\text{\tiny{CC}}}$ is the weight, ${\bf{W}}$, that minimizes the asymptotic MSE of estimators of the form $\hat{\bm{\beta}}^{\text{\tiny{lab}}}-{{\bf{W}}}\left(\hat{\bm{\gamma}}^{\text{\tiny{lab}}}-\hat{\bm{\gamma}}^{\text{\tiny{all}}}\right)$ for estimating $\bm{\beta}^{*}$.
Proof  Letting $\hat{\bm{\beta}}_{{\bf{W}}}=\hat{\bm{\beta}}^{\text{\tiny{lab}}}-{\bf{W}}(\hat{\bm{\gamma}}^{\text{\tiny{lab}}}-\hat{\bm{\gamma}}^{\text{\tiny{all}}})$, the MSE of estimating $\hat{\bm{\beta}}_{{\bf{W}}}$ is

### v3

@@ B.1 Asymptotic expansion of $\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$
Theorem S1 (Influence function expansion of $\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$).
@@
Proof.
The only difference in studying the asymptotic properties of $\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$ compared to $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$ is that we rely on the asymptotic linear expansion of $\hat{\bm{\gamma}}^{\text{\tiny{unlab}}}$ rather than $\hat{\bm{\gamma}}^{\text{\tiny{all}}}$. Using the same argument as in the proof of Theorem 1,
@@ B.2 Comparison of $\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$ $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$
Corollary S1.1 (Comparison of $\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$ with $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$).
@@
$\Delta_{\text{aVar}}(\hat{\bm{\beta}}^{\text{\tiny{lab}}},\hat{\bm{\beta}}^{\text{\tiny{PPI}}})$ $=\frac{1}{\pi}{\bf{A}}\left(2\bm{\Sigma}_{Y,\widehat{Y}}-\frac{1}{1-\pi}\bm{\Sigma}_{\widehat{Y},\widehat{Y}}\right){\bf{A}}.$
$\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$ is therefore more efficient than $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ if and only if $2\bm{\Sigma}_{Y,\widehat{Y}}\succ\frac{1}{1-\pi}\bm{\Sigma}_{\widehat{Y},\widehat{Y}}$. A sufficient condition for $\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$ to have lower variance than $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ is that
@@
Proof.
From the proof for Corollary 1.1, $n[\text{aVar}(\hat{\bm{\beta}}^{\text{\tiny{lab}}})]=\frac{1}{\pi}{\bf{A}}\bm{\Sigma}_{Y,Y}{\bf{A}}.$ By Theorem S1,
@@
$=\frac{1}{\pi}{\bf{A}}\bm{\Sigma}_{Y,Y}{\bf{A}}+\frac{1}{\pi(1-\pi)}{\bf{A}}\bm{\Sigma}_{\widehat{Y},\widehat{Y}}{\bf{A}}-\frac{2}{\pi}{\bf{A}}\bm{\Sigma}_{Y,\widehat{Y}}{\bf{A}}$
$=\frac{1}{\pi}{\bf{A}}\left[\bm{\Sigma}_{Y,Y}+\frac{1}{1-\pi}\bm{\Sigma}_{\widehat{Y},\widehat{Y}}-2\bm{\Sigma}_{Y,\widehat{Y}}\right]{\bf{A}}.$
@@
$\Delta_{\text{aVar}}(\hat{\bm{\beta}}^{\text{\tiny{lab}}},\hat{\bm{\beta}}^{\text{\tiny{PPI}}})=\frac{1}{\pi}{\bf{A}}\left[2\bm{\Sigma}_{Y,\widehat{Y}}-\frac{1}{1-\pi}\bm{\Sigma}_{\widehat{Y},\widehat{Y}}\right]{\bf{A}}.$
$\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$ is therefore more efficient than $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ if and only if $2\bm{\Sigma}_{Y,\widehat{Y}}\succ\frac{1}{1-\pi}\bm{\Sigma}_{\widehat{Y},\widehat{Y}}$. Note that if $2\mathbb{E}\left(\epsilon_{Y}\epsilon_{\widehat{Y}}\mid{\bf{X}}\right)>\frac{1}{1-\pi}\mathbb{E}\left(\epsilon_{\widehat{Y}}^{2}\mid{\bf{X}}\right)$ or equivalently if $\mathbb{E}\left[\left(\epsilon_{\widehat{Y}}-\epsilon_{Y}\right)^{2}\mid{\bf{X}}\right]<\mathbb{E}\left(\epsilon_{Y}^{2}\mid{\bf{X}}\right)+\left(1-\frac{1}{1-\pi}\right)\mathbb{E}\left(\epsilon_{\widehat{Y}}^{2}\mid{\bf{X}}\right),$ the diagonal entries of $\Delta_{\text{aVar}}(\hat{\bm{\beta}}^{\text{\tiny{lab}}},\hat{\bm{\beta}}^{\text{\tiny{PPI}}})$ are positive and $\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$ will have a lower variance than $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$. ∎
@@ B.3 Comparison of $\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$ $\hat{\bm{\beta}}^{\text{\tiny{PPI}}_{a}}$
$2\bm{\Sigma}_{Y,\widehat{Y}}\succ\frac{1}{1-\pi}\bm{\Sigma}_{\widehat{Y},\widehat{Y}}$
@@
$2\bm{\Sigma}_{Y,\widehat{Y}}\succ\bm{\Sigma}_{\widehat{Y},\widehat{Y}}.$
@@
$=\left(\frac{1}{\pi}-1\right){\bf{A}}\left(2\bm{\Sigma}_{Y,\widehat{Y}}-\bm{\Sigma}_{\widehat{Y},\widehat{Y}}\right){\bf{A}}-\frac{1}{\pi}{\bf{A}}\left(2\bm{\Sigma}_{Y,\widehat{Y}}-\frac{1}{1-\pi}\bm{\Sigma}_{\widehat{Y},\widehat{Y}}\right){\bf{A}}$
$={\bf{A}}\left(\frac{2-\pi}{1-\pi}\bm{\Sigma}_{\widehat{Y},\widehat{Y}}-2\bm{\Sigma}_{Y,\widehat{Y}}\right){\bf{A}}.$
Note that if $\hat{\bm{\beta}}^{\text{\tiny{PPI}}}$ is less efficient than $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$ then $\frac{1}{1-\pi}\bm{\Sigma}_{\widehat{Y},\widehat{Y}}\succ 2\bm{\Sigma}_{Y,\widehat{Y}}$, implying
$\frac{2-\pi}{1-\pi}\bm{\Sigma}_{\widehat{Y},\widehat{Y}}\succ 2\bm{\Sigma}_{Y,\widehat{Y}}.$
@@ C.1 Proof of Theorem 2
Theorem 2 (Influence function expansion of $\hat{\bm{\beta}}^{\text{\tiny{CC}}}$).
Following Chen and Chen [2000],
@@
Proof.
The proof follows from Chen and Chen [2000], but is detailed for completeness. Recall that
$\begin{split}\widehat{{\bf{W}}}^{\text{\tiny{CC}}}&=\widehat{{\bf{A}}}\left\{\widehat{\text{Cov}}\left[\phi(Y,{\bf{X}};\bm{\beta}^{*}),\phi(\widehat{Y},{\bf{X}};\bm{\gamma}^{*})\right]\right\}\{\widehat{\text{Var}}[\phi(\widehat{Y},{\bf{X}};\bm{\gamma}^{*})]\}^{-1}\widehat{{\bf{A}}}^{-1}\\ &=\widehat{{\bf{A}}}\widehat{\bm{\Sigma}}_{Y,\widehat{Y}}\widehat{\bm{\Sigma}}^{-1}_{\widehat{Y},\widehat{Y}}\widehat{{\bf{A}}}^{-1}\end{split}$
@@
$\begin{split}&\widehat{{\bf{A}}}=n_{\text{\tiny{lab}}}\left(\sum_{i=1}^{n_{\text{\tiny{lab}}}}{\bf{X}}_{i}{\bf{X}}_{i}^{{\sf\scriptscriptstyle{T}}}\right)^{-1},\\ &\widehat{\bm{\Sigma}}_{Y,\widehat{Y}}=n_{\text{\tiny{lab}}}^{-1}\sum_{i=1}^{n_{\text{\tiny{lab}}}}\left(Y_{i}-{\bf{X}}_{i}^{{\sf\scriptscriptstyle{T}}}\hat{\bm{\beta}}^{\text{\tiny{lab}}}\right)\left(\hat{Y}_{i}-{\bf{X}}_{i}^{{\sf\scriptscriptstyle{T}}}\hat{\bm{\gamma}}^{\text{\tiny{lab}}}\right){\bf{X}}_{i}{\bf{X}}_{i}^{{\sf\scriptscriptstyle{T}}},\\ &\mbox{and }\widehat{\bm{\Sigma}}_{\widehat{Y},\widehat{Y}}=n_{\text{\tiny{lab}}}^{-1}\sum_{i=1}^{n_{\text{\tiny{lab}}}}\left(\hat{Y}_{i}-{\bf{X}}_{i}^{{\sf\scriptscriptstyle{T}}}\hat{\bm{\gamma}}^{\text{\tiny{lab}}}\right)^{2}{\bf{X}}_{i}{\bf{X}}_{i}^{{\sf\scriptscriptstyle{T}}}.\end{split}$
@@
$n^{\frac{1}{2}}\left[\widehat{{\bf{W}}}^{\text{\tiny{CC}}}(\hat{\bm{\gamma}}^{\text{\tiny{lab}}}-\bm{\gamma}^{*})\right]={\bf{A}}\bm{\Sigma}_{Y,\widehat{Y}}\bm{\Sigma}^{-1}_{\widehat{Y},\widehat{Y}}\left\{n^{-\frac{1}{2}}\sum_{i=1}^{n}\left[{\bf{X}}_{i}\left(\widehat{Y}_{i}-{\bf{X}}_{i}^{{\sf\scriptscriptstyle{T}}}\bm{\gamma}^{*}\right)\right]\frac{R_{i}}{\pi}\right\}+o_{p}(1)$
@@
$n^{\frac{1}{2}}\left[\widehat{{\bf{W}}}^{\text{\tiny{CC}}}(\hat{\bm{\gamma}}^{\text{\tiny{all}}}-\bm{\gamma}^{*})\right]={\bf{A}}\bm{\Sigma}_{Y,\widehat{Y}}\bm{\Sigma}^{-1}_{\widehat{Y},\widehat{Y}}\left\{n^{-\frac{1}{2}}\sum_{i=1}^{n}\left[{\bf{X}}_{i}\left(\widehat{Y}_{i}-{\bf{X}}_{i}^{{\sf\scriptscriptstyle{T}}}\bm{\gamma}^{*}\right)\right]\right\}+o_{p}(1).$
@@
Corollary 2.1 ($\hat{\bm{\beta}}^{\text{\tiny{CC}}}$ is at least as efficient as $\hat{\bm{\beta}}^{\text{\tiny{lab}}}$).
@@
$\Delta_{\text{aVar}}(\hat{\bm{\beta}}^{\text{\tiny{lab}}},\hat{\bm{\beta}}^{\text{\tiny{CC}}})=\left(\frac{1}{\pi}-1\right){\bf{A}}\bm{\Sigma}_{Y,\widehat{Y}}\bm{\Sigma}^{-1}_{\widehat{Y},\widehat{Y}}\bm{\Sigma}_{Y,\widehat{Y}}{\bf{A}}\succeq{\bf{0}}.$
@@
Proof.
From the proof for Corollary 1.1, $n[\text{aVar}(\hat{\bm{\beta}}^{\text{\tiny{lab}}})]=\frac{1}{\pi}{\bf{A}}\bm{\Sigma}_{Y,Y}{\bf{A}}.$ By Theorem 2.1, $n[\text{aVar}(\hat{\bm{\beta}}^{\text{\tiny{CC}}})]$ is given by
$\begin{split}&\mathbb{E}\left\{\left[{\bf{A}}\bm{\phi}(Y,{\bf{X}};\bm{\beta}^{*})\right]^{\otimes 2}\frac{R^{2}}{\pi^{2}}\right\}+\mathbb{E}\left\{\left[{\bf{A}}\bm{\Sigma}_{Y,\widehat{Y}}\bm{\Sigma}^{-1}_{\widehat{Y},\widehat{Y}}\bm{\phi}(\widehat{Y},{\bf{X}};\bm{\gamma}^{*})\right]^{\otimes 2}\frac{(\pi-R)^{2}}{\pi^{2}}\right\}\\ +&2\mathbb{E}\left\{\left[{\bf{A}}\bm{\phi}(Y,{\bf{X}};\bm{\beta}^{*})\right]\left[{\bf{A}}\bm{\Sigma}_{Y,\widehat{Y}}\bm{\Sigma}^{-1}_{\widehat{Y},\widehat{Y}}\bm{\phi}(\widehat{Y},{\bf{X}};\bm{\gamma}^{*})\right]^{{\sf\scriptscriptstyle{T}}}\frac{R(\pi-R)}{\pi^{2}}\right\}\\ =&\frac{1}{\pi}{\bf{A}}\bm{\Sigma}_{Y,Y}{\bf{A}}+\frac{1-\pi}{\pi}{\bf{A}}\bm{\Sigma}_{Y,\widehat{Y}}\bm{\Sigma}^{-1}_{\widehat{Y},\widehat{Y}}\bm{\Sigma}_{Y,\widehat{Y}}{\bf{A}}-2\frac{1-\pi}{\pi}{\bf{A}}\bm{\Sigma}_{Y,\widehat{Y}}\bm{\Sigma}^{-1}_{\widehat{Y},\widehat{Y}}\bm{\Sigma}_{Y,\widehat{Y}}{\bf{A}}\\ =&\frac{1}{\pi}{\bf{A}}\bm{\Sigma}_{Y,Y}{\bf{A}}-\frac{1-\pi}{\pi}{\bf{A}}\bm{\Sigma}_{Y,\widehat{Y}}\bm{\Sigma}^{-1}_{\widehat{Y},\widehat{Y}}\bm{\Sigma}_{Y,\widehat{Y}}{\bf{A}}.\end{split}$
Hence, $\Delta_{\text{aVar}}(\hat{\bm{\beta}}^{\text{\tiny{lab}}},\hat{\bm{\beta}}^{\text{\tiny{CC}}})=\frac{1-\pi}{\pi}{\bf{A}}\bm{\Sigma}_{Y,\widehat{Y}}\bm{\Sigma}^{-1}_{\widehat{Y},\widehat{Y}}\bm{\Sigma}_{Y,\widehat{Y}}{\bf{A}}\succeq 0$. ∎
@@
Corollary 2.2 ($\hat{\bm{\beta}}^{\text{\tiny{CC}}}$ is the most efficient estimator in its class).
Let the mean squared error (MSE) of $\hat{\bm{\beta}}$ for estimating $\bm{\beta}^{*}$ be defined as $\mathbb{E}[(\hat{\bm{\beta}}-\bm{\beta}^{*})^{{\sf\scriptscriptstyle{T}}}(\hat{\bm{\beta}}-\bm{\beta}^{*})]$. Analogously to the well-known result for the best linear predictor (e.g., Bickel and Doksum [2015], page 40), ${\bf{W}}^{\text{\tiny{CC}}}$ is the weight, ${\bf{W}}$, that minimizes the asymptotic MSE of estimators of the form $\hat{\bm{\beta}}^{\text{\tiny{lab}}}-{{\bf{W}}}\left(\hat{\bm{\gamma}}^{\text{\tiny{lab}}}-\hat{\bm{\gamma}}^{\text{\tiny{all}}}\right)$ for estimating $\bm{\beta}^{*}$.
Proof.
Letting $\hat{\bm{\beta}}_{{\bf{W}}}=\hat{\bm{\beta}}^{\text{\tiny{lab}}}-{\bf{W}}(\hat{\bm{\gamma}}^{\text{\tiny{lab}}}-\hat{\bm{\gamma}}^{\text{\tiny{all}}})$, the MSE of estimating $\hat{\bm{\beta}}_{{\bf{W}}}$ is